xx IDBinBH

IDxBinxBH

Erik Schnetter <schnetter@cct.lsu.edu>

November 24 2009

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11filename 10Scope: privatex01 STRING

11Description: Input file name

11 Range 10  x01  Default: (none)

11           10

11

11initialxdata 10Scope: shared from ADMBASEx01  KEYWORD

11Extends ranges:

11IDxBinxBH 10

11

11initialxdtlapse 10Scope: shared from ADMBASEx01  KEYWORD

11Extends ranges:

11IDxBinxBH 10

11

11initialxdtshift10Scope: shared from ADMBASEx01  KEYWORD

11Extends ranges:

11IDxBinxBH 10

11

11initialxlapse 10Scope: shared from ADMBASEx01  KEYWORD

11Extends ranges:

11IDxBinxBH 10

11

11initialxshift 10Scope: shared from ADMBASEx01  KEYWORD

11Extends ranges:

11IDxBinxBH 10

11

7 Interfaces

General Implements: idxbinxbh

Inherits: grid summationbyparts admbase 8 Schedule This section lists all the variables which are assigned storage by thorn EinsteinInitialData/MeudonxBinxBH. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxPARAMCHECK (conditional) x idxbinxbhxcheckxparameters x check parameters x

11x11Language:11c
11 x 11Type: 11function

11

ADMBasexInitialData (conditional) x idxbinxbhxinitialise x set up binary black hole initial data x

11x11Language:11c
11 x 11Reads: 11grid::coordinates(everywhere)
11 x 11Type: 11function
11 x 11Writes: 11admbase::metric(everywhere)
11 x 11x  11admbase::shift(everywhere)
11 x 11x  11admbase::dtshift(everywhere)
11 x 11x  11admbase::curv(everywhere)
11 x 11x  11admbase::alp(everywhere)
11 x 11x  11admbase::dtalp(everywhere)